OpenStudy (anonymous):

Solve x2 + 4x – 12 = 0 by completing the square. Show your work for full credit.

OpenStudy (espex):

When completing the square you want to start by putting your equation in the form of \(ax^2+bx=c\) You then take \(\frac{b}{2}\), square it, and add it to both sides of the equals sign \(ax^2+bx+d=(c+d)\). Looking at the left hand side only, you will notice that your first and last terms are perfect squares, convert the left to squared form \((\sqrt{a}+\sqrt{d})^2\). Take the square root of both sides and solve for 'x'.

OpenStudy (anonymous):

@eSpeX thank you so much